# Fractionally modulated discrete Carleson’s Theorem and pointwise Ergodic Theorems along certain curves

### v1

@@ 6. Concluding the Proof of Theorem 1
$\begin{split}&\sup_{\lambda\in[-1/2,1/2)}\Big{|}\sum_{j\leq 0}\int_{\mathbb{R}}(f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}])(x-t)\big{(}e(\lambda[t]_{\mathsf{i}}^{c})-1\big{)}\psi_{j}(t){\rm d}t\Big{|}\\ &\lesssim\sup_{\lambda\in[-1/2,1/2)}\sum_{j\leq 0}\int_{\mathbb{R}}|f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]|(x-t)\lambda|t|^{c}|\psi_{j}(t)|{\rm d}t\\ &\lesssim\sum_{j\leq 0}2^{jc}\int_{\mathbb{R}}|f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]|(x-t)|\psi_{j}(t)|{\rm d}t\leq\sum_{j\leq 0}2^{jc}M\big{(}f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]\big{)}(x)\lesssim M\big{(}f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]\big{)}(x)\text{,}\end{split}$

### v2

@@ 6. Concluding the Proof of Theorem 1
$\begin{split}&\sup_{\lambda\in[-1/2,1/2)}\Big|\sum_{j\leq 0}\int_{\mathbb{R}}(f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}])(x-t)\big(e(\lambda[t]_{\mathsf{i}}^{c})-1\big)\psi_{j}(t){\rm d}t\Big|\\ &\lesssim\sup_{\lambda\in[-1/2,1/2)}\sum_{j\leq 0}\int_{\mathbb{R}}|f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]|(x-t)|\lambda||t|^{c}|\psi_{j}(t)|{\rm d}t\\ &\lesssim\sum_{j\leq 0}2^{jc}\int_{\mathbb{R}}|f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]|(x-t)|\psi_{j}(t)|{\rm d}t\leq\sum_{j\leq 0}2^{jc}M\big(f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]\big)(x)\lesssim M\big(f*\mathcal{F}_{\mathbb{R}}^{-1}[\eta_{1}]\big)(x)\text{,}\end{split}$
@@
Let us mention that by combining Lemma 6.2, Lemma 6.3, as well as the fact that our handling of $S_{3}$ clearly yields constants independent of $c$ we get
(6.16) $\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\mathcal{F}^{-1}_{\mathbb{R}}\Big[\sum_{j\in\mathbb{N}}\mathcal{F}_{\mathbb{R}}[f](\cdot)H_{j}^{\mathsf{i}}(\cdot,\lambda)\bm{1}_{X_{j}}(\lambda)\Big]\Big|\bigg\|_{L^{2}(\mathbb{R})}\lesssim(c-1)^{-\frac{1}{2}}\|f\|_{L^{2}(\mathbb{R})}\text{,}$
and returning to (5.12), we get that
(6.17) $\sup_{f\in\ell^{2}:\,\|f\|_{\ell^{2}}\leq 1}\bigg\|\sup_{\lambda\in[-1/2,1/2)}\Big|\sum_{n\neq 0}f(x-n)\frac{e(\lambda\mathsf{sign}^{\mathsf{i}}(n)\lfloor|n|^{c}\rfloor)}{n}\Big|\bigg\|_{{\ell_{{\rm d}x}^{2}}}\lesssim(c-1)^{-\frac{1}{2}}(2-c)^{-\frac{13}{4}}\text{,}\quad\text{as claimed earlier.}$